OpenStudy (anonymous):

The instructions for baking a turkey recommend cooking a 12-pound bird for 30 minutes in a 350 degree oven. The instructions also suggest adding 3 extra minutes of cooking time for each additional pound. The cooking time for a turkey can be calculated using the equation below, where w represents the wieght of the turkey. If Maria needs to cook her turkey for 45 minutes, how many pounds does Maria's turkey weigh? Cooking Time = 3(w - 12) + 30

OpenStudy (anonymous):

45=3(w-12)+30 SOlve for w 45=3w-36+30 combine like terms 45=3w-6 add 6 to both sides 51=3w divide each side by 3 w=17lbs